OpenStudy (anonymous):

Solve x3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

You have\[x^3 = \frac{ 1 }{ 8 }\] If you can express \(\frac{1}{8}\) as a number raised to exponent 3, then you can simply equate the bases. Can you think of a number that, when cubed, equals 1\8 ?

OpenStudy (anonymous):

1/2 @ospreytriple

OpenStudy (anonymous):

b is incorrect because the plus and minus sign you dont use in power to the third

OpenStudy (anonymous):

Excellent. So\[x^3 = \left( \frac{ 1 }{ 2 } \right)^3\]When the exponents are the same, simply equate the bases. So\[x=?\]

OpenStudy (anonymous):

1/2 is the answer its not 1/2 plus minus sign

OpenStudy (anonymous):

@Marc1313 , please do not provide direct answers. It does not help the asker and it violates OpenStudy's Code of Conduct.

OpenStudy (anonymous):

i didnt provide direct answer @GIL.ojei did i was then one who said 1/2 plus minus sign so i was jsust saying that that is incorrect so idk what your talking about

OpenStudy (anonymous):

So sorry, @Marc1313 . My mistake, I thought you were a third party for a second. Your answer is correct. Well done.

OpenStudy (anonymous):

Okay. thank you.

OpenStudy (anonymous):

wait which one is the answer @ospreytriple

OpenStudy (anonymous):

We worked it down to\[x^3 = \left( \frac{ 1 }{ 2 } \right)^3\]Since exponents are the same we can equate just the bases, i.e.\[x=\frac{ 1 }{ 2 }\]See how that works?